OpenStudy (anonymous):

If B is between A and C find the value of BC. AB = 3, BC = 5x - 1, AC = 12

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

if B between A & C (assume on the same line), then: 3+(5x-1)=12 5x=10 x=2

OpenStudy (anonymous):

since x=2, so, BC=5*2-1=9